A local sports bar routinely promotes sporting events, such as the Super Bowl, the NCAA Sweet Sixteen tournament, the BCS games, the Masters, NBA and NFL playoffs, NHL, baseball, NASCAR, and key curling bonspiels. By keeping records, the owner has determined that when the cover price is $13 the average number of patrons is 214. For every $1 change in cover charge, the number of patrons changes by 12. Assuming a linear demand curve, calculate the maximum willingness to buy for sporting events at this sports bar.
Answer #1

The equation for demand can be found as
Given Slope is 12
Intercept is 214
P0 is 13
so putting in the equation
Demand=Intercept-Slope*(P-P0)
Q=214-12*(P-13)

Maximum willingness occurs for Q=1
Hence,
P=213/12+13=30.75